Does the point (6, 9) satisfy the inequality y  8x – 2? (y is less than or equal to 8x-2)
Answer by london maths tutor(243)

Put x = 6 and y = 9
8*6 - 2 = 48 -2 = 46 which is bigger than 9
Therefore, y < 8x - 6